# Scalable and Accurate Graph Reasoning with LLM-based Multi-Agents

### v2

@@ 5.3 Experiment 3: Case Study
Although GraphWiz performed well on fine-tuned tasks, it exhibits severe overfitting when faced with real-world graph problems, failing to apply the graph reasoning knowledge learned during the fine-tuning phase. Since GraphWiz uses a consistent graph node description, the sentence "The nodes are numbered from 0 to …" appears across all datasets during the mixed-task instruction tuning. When the actual problem has nodes numbered from 1 to 20, it still assumes the existence of node 0. As a result, both GraphWiz models first output that the graph has 21 nodes and an incorrect number of edges. Furthermore, neither of the two GraphWiz models recognizes that this is a problem associated with web page importance ranking. Instead, they approach it as the bipartite graph check or topological sort problems they had been fine-tuned on. Additionally, neither model generates an explicit and correct reasoning path. These observations indicate that there is still a significant gap between excelling in classic graph reasoning tasks and effectively solving real-world graph reasoning problems. In contrast, GAR correctly identifies that the problem should be solved using knowledge related to PageRank (yang2024pagerank) and designs an algorithm that adhered to the distributed paradigm (Note: the distributed algorithm library does not contain a PageRank algorithm template). GAR then assigns the algorithm to agent nodes for execution, ultimately obtaining the PageRank value for each node and arriving at the correct conclusion. Through the distributed paradigm, GAR effectively bridges the powerful knowledge learned by LLMs with the solving of real-world graph reasoning problems, which enables it to flexibly handle practical issues in a distributed manner. This case study demonstrates the feasibility of using GAR to solve real-world graph reasoning problems, indicating its substantial practical applicability and offering researchers and practitioners a powerful framework to address such tasks.
@@ 6 Conclusion
Appendix A Distributed algorithms under the distributed paradigm
A.1 Example of distributed algorithms in distributed algorithm library
Shortest Path: See Figure 6.
Connectivity: See Figure 7.
A.2 Example of distributed algorithms designed by the Master LLM
PageRank: See Figure 8.
Hamilton Path: See Figure 9.
Subgraph Matching: See Figure 10.
Appendix B The GraphInstruct Dataset
The statistics and detailed information of GraphInstruct are shown in Table 3. Hamilton Path and Subgraph Matching are NP-complete problems.
@@ Appendix B The GraphInstruct Dataset
### State
1. ‘visited‘: Boolean indicating if the node has been visited.
2. ‘path_length‘: Integer representing the current length of the path.
3. ‘max_path_length‘: Integer tracking the longest path found.
### Initialization
Step 1: Set ‘visited = False‘, ‘path_length = 0‘, and ‘max_path_length = 1‘ for all nodes.
Step 2: For the initial node (e.g., Node Id: 0), set ‘visited = True‘ and
‘path_length = 1‘.
Distributed_Algorithm_Designer: ### Message
1. ‘path_length‘: The current path length from the sender.
2. ‘max_path_length‘: The maximum path length known to the sender.
3. ‘visited_flag‘: Boolean indicating whether the sender has visited the node.
### Send
Step 1: For each neighbor (Node A, Node B), if ‘visited‘ is ‘False‘, construct the message as follows:
- Message to Node A: ‘(Node Id: <node_id>, path_length + 1, max(max_path_length, path_length + 1))‘.
- Message to Node B: ‘(Node Id: <node_id>, path_length + 1, max(max_path_length, path_length + 1))‘.
Step 2: Send the constructed messages to both neighbors.
### Example
## Input
Node Id: 0
State:
1. visited: False
2. path_length: 2
3. max_path_length: 3
Neighbor Information:
Connected to:
Node A
Node B
## Process
Step 1: Check neighbors Node A and Node B. Since ‘visited‘ is
‘False‘, proceed to construct messages.
- For Node A:
- Calculate new ‘path_length‘: ‘2 + 1 = 3‘
- Calculate ‘max_path_length‘: ‘max(3, 2 + 1) = 3‘
- For Node B:
- Same calculations apply as for Node A.
Step 2: Construct messages:
- Message to Node A: ‘(Node Id: 0, 3, 3)‘
- Message to Node B: ‘(Node Id: 0, 3, 3)‘
## Output
Message sent to Node A:
1. path_length: 3
2. max_path_length: 3
3. visited_flag: False
Message sent to Node B:
1. path_length: 3
2. max_path_length: 3
3. visited_flag: False
### Update
Step 1: For each received message, check the ‘visited_flag‘. If it is
‘False‘, proceed to the next steps.
Step 2: Update ‘visited‘ to ‘True‘ for the current node.
Step 3: Update ‘path_length‘ to the maximum of its current value and
the received ‘path_length‘ from the message, incrementing by 1 if the
message indicates a potential continuation of the path.
Step 4: Update ‘max_path_length‘ to the maximum of its current value
and the received ‘max_path_length‘ from the message.
Step 5: If any messages were processed, trigger a re-evaluation of
outgoing messages to neighbors, if applicable.
### Example
## Input
Node Id: 1
State:
1. visited: False
2. path_length: 1
3. max_path_length: 2
Received Messages:
Message 1:
1. path_length: 2
2. max_path_length: 3
3. visited_flag: False
Message 2:
1. path_length: 3
2. max_path_length: 4
3. visited_flag: True
## Process
Step 1: Process Message 1. Since ‘visited_flag‘ is ‘False‘, proceed:
- Update ‘visited‘ to ‘True‘.
- Update ‘path_length‘: ‘max(1, 2) + 1 = 3‘.
- Update ‘max_path_length‘: ‘max(2, 3) = 3‘.
Step 2: Process Message 2. Since ‘visited_flag‘ is ‘True‘, ignore this
message and do not make any changes.
## Output
State:
1. visited: True
2. path_length: 3
3. max_path_length: 3
### Termination
The algorithm continues until:
- A node reaches a ‘path_length‘ equal to the total number of nodes
(indicating a Hamiltonian path exists).
- Or until a predefined number of iterations equal to the total number
- of nodes (‘N‘) is reached without discovering a Hamiltonian path.
Master: ##################### ITER 3, STEP 1: REASONING ######################
Master: {’thought’: "The algorithm is designed to check for a
Hamiltonian path. Now let’s execute it to see if such a path exists.",
’speak’: "The algorithm is ready. Let’s execute it to determine if there is
a Hamiltonian path.", ’function’: [{’name’: ’Distributed_Execution’,
’arguments’: {}}]}
Master: ####################### ITER 3, STEP 2: ACTING #######################
0: State: 1. visited: True 2. path_length: 1 3. max_path_length: 1
2: State: 1. visited: False 2. path_length: 0 3. max_path_length: 1
5: State: 1. visited: False 2. path_length: 0 3. max_path_length: 1
3: State: 1. visited: False 2. path_length: 0 3. max_path_length: 1
4: State: 1. visited: False 2. path_length: 0 3. max_path_length: 1
1: State: 1. visited: False 2. path_length: 0 3. max_path_length: 1
Node 5 Send Message to Node 1: 1. path_length: 1 2. max_path_length:
1 3. visited_flag: False
Node 5 Send Message to Node 2: 1. path_length: 1 2. max_path_length:
1 3. visited_flag: False
Node 5 Send Message to Node 3: 1. path_length: 1 2. max_path_length:
1 3. visited_flag: False
All agents’ state unchanged, terminating early...
>>> END
system: 1. Execute function Distributed_Execution
[ARGUMENTS]:
[STATUS]: SUCCESS
[RESULT]: Node: 0 State: 1. visited: True 2. path_length: 5 3. max_path_length: 1
Node: 1 State: 1. visited: True 2. path_length: 2 3. max_path_length: 1
Node: 2 State: 1. visited: True 2. path_length: 4 3. max_path_length: 3
Node: 3 State: 1. visited: True 2. path_length: 5 3. max_path_length: 3
Node: 4 State: 1. visited: True 2. path_length: 4 3. max_path_length: 3
Node: 5 State: 1. visited: True 2. path_length: 5 3. max_path_length: 3
Master: ##################### ITER 4, STEP 1: REASONING ######################
Master: {’thought’: ’Since the maximum path length found is 5, which is
less than the total number of nodes (6), there is no Hamiltonian path in
the graph.’, ’speak’: ’No, there is no Hamiltonian path in this graph.’,
’function’: []}
Appendix C Execution examples of GraphAgent-Reasoner
@@
Master: ##################### ITER 1, STEP 1: REASONING ######################
Master: {’thought’: ’I need to find the shortest distances from node 1 to all
other nodes in the given undirected graph. First, I will build the graph using
the provided edges and then design a distributed algorithm to find the
shortest paths.’, ’speak’: ’Building the graph and designing the distributed
algorithm to find the shortest distances from node 1.’, ’function’: ’[{"name":
"Build_Graph", "arguments": {"Edges": [[0, 7, 9], [0, 1, 7], [0, 4, 9], [1, 7, 1], [2,
7, 7], [2, 6, 5], [2, 5, 8], [3, 5, 9], [3, 4, 8], [3, 6, 1], [4, 7, 7], [4, 5, 6], [5, 7, 6]],
"Nodes_Num": 8, "Directed": false}}]’}
Master: ####################### ITER 1, STEP 2: ACTING #######################
>>> Edges: [[0, 7, 9], [0, 1, 7], [0, 4, 9], [1, 7, 1], [2, 7...
>>> Nodes_Num: 8
>>> Directed: False
>>> END
system: 1. Execute function Build_Graph
[ARGUMENTS]:
Edges: [[0, 7, 9], [0, 1, 7], [0, 4, 9], [1, 7, 1], [2, 7, 7], [2, 6, 5], [2, 5, 8], [3,
5, 9], [3, 4, 8], [3, 6, 1], [4, 7, 7], [4, 5, 6], [5, 7, 6]]
Nodes_Num: 8
Directed: False
[STATUS]: SUCCESS
Master: ##################### ITER 2, STEP 1: REASONING ######################
Master: {’thought’: ’The graph has been successfully built. Now, I will design
a distributed algorithm to find the shortest distances from node 1 to all
other nodes.’, ’speak’: ’Designing a distributed algorithm to find the
shortest distances from node 1.’, ’function’: ’[{"name":
"Design_Distributed_Algorithm", "arguments": {"Problem_Str": "Find the
shortest distance from node 1 to all other nodes in an undirected graph
with 8 nodes and given edges."}}]’}
Master: ####################### ITER 2, STEP 2: ACTING #######################
>>> END
Master: ##################### ITER 3, STEP 1: REASONING ######################
Master: {’thought’: ’The distributed algorithm has been successfully
designed. Now, I will execute the algorithm to find the shortest distances
from node 1 to all other nodes.’, ’speak’: ’Executing the distributed
algorithm to calculate the shortest distances from node 1.’, ’function’:
’[{"name": "Distributed_Execution", "arguments": {}}]’}
Master: ####################### ITER 3, STEP 2: ACTING #######################
Initialization:
1: State: 1. distance: 0
[…]
Node 1 Send Message to Node 0: 1. new_distance: 7
Node 1 Send Message to Node 7: 1. new_distance: 1
[…]
All agents’ state unchanged, terminating early...
>>> END
system: 1. Execute function Distributed_Execution
[ARGUMENTS]:
[STATUS]: SUCCESS
[RESULT]: Node: 0 State: 1. distance: 7
Node: 1 State: 1. distance: 0
Node: 2 State: 1. distance: 8
Node: 3 State: 1. distance: 14
Node: 4 State: 1. distance: 8
Node: 5 State: 1. distance: 7
Node: 6 State: 1. distance: 13
Node: 7 State: 1. distance: 1
Master: ##################### ITER 4, STEP 1: REASONING ######################
The shortest distances from node 1 are as follows: Node 0: 7, Node 1: 0,
Node 2: 8, Node 3: 14, Node 4: 8, Node 5: 7, Node 6: 13, Node 7: 1.Master:
{’thought’: ’The distributed algorithm has been successfully executed, and I
have obtained the shortest distances from node 1 to all other nodes.’,
’speak’: ’The shortest distances from node 1 are as follows: Node 0: 7, Node
1: 0, Node 2: 8, Node 3: 14, Node 4: 8, Node 5: 7, Node 6: 13, Node 7: 1.’,
’function’: []}

### v3

@@ 5.3 Case Study
Although GraphWiz performed well on fine-tuned tasks, it exhibits severe overfitting when faced with real-world graph problems, failing to apply the graph reasoning knowledge learned during the fine-tuning phase. Since GraphWiz uses a consistent graph node description, the sentence ”The nodes are numbered from 0 to …” appears across all datasets during the multi-task instruction tuning. When the actual problem has nodes numbered from 1 to 20, it still assumes the existence of node 0. Both GraphWiz models output that the graph has 21 nodes with an incorrect number of edges.
Furthermore, neither of the two GraphWiz models recognizes that this is a problem associated with web page importance ranking. Instead, they approach it as the bipartite graph check or topological sort problems they had been fine-tuned on. Additionally, neither model generates an explicit and correct reasoning path. These observations indicate that there is still a significant gap between excelling in classic graph reasoning tasks and effectively solving real-world graph reasoning problems. In contrast, GAR correctly identifies that the problem should be solved using knowledge related to PageRank (yang2024pagerank) and designs an algorithm that adhered to the distributed paradigm (Note: the distributed algorithm library does not contain a PageRank algorithm template). GAR then assigns the algorithm to agent nodes for execution, ultimately obtaining the PageRank value for each node and arriving at the correct conclusion. Through the distributed paradigm, GAR effectively bridges the powerful knowledge learned by LLMs with the solving of real-world graph reasoning problems, which enables it to flexibly handle practical issues in a distributed manner. This case study demonstrates the feasibility of using GAR to solve real-world graph reasoning problems, indicating its substantial practical applicability and offering researchers and practitioners a powerful framework to address such tasks.
@@ 6 Conclusion
7 Acknowledgments
This work was supported in part by Ant Group through Ant Research Intern Program, by National Natural Science Foundation of China (No. 92470128, No. U2241212), by National Science and Technology Major Project (2022ZD0114802), by Beijing Outstanding Young Scientist Program No.BJJWZYJH012019100020098, by the National Key Research and Development Plan of China (2023YFB4502305). We also wish to acknowledge the support provided by the fund for building world-class universities (disciplines) of Renmin University of China, by Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, by Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Public Policy and Decision-making Research Lab, and Public Computing Cloud, Renmin University of China. The work was partially done at Gaoling School of Artificial Intelligence, Beijing Key Laboratory of Big Data Management and Analysis Methods, MOE Key Lab of Data Engineering and Knowledge Engineering, and Pazhou Laboratory (Huangpu), Guangzhou, Guangdong 510555, China.
8 Appendix
8.1 The GraphInstruct Dataset
The statistics and detailed information of GraphInstruct are shown in Table 3. Hamilton Path and Subgraph Matching are NP-complete problems. However, there are no well-established distributed algorithms for these two problems, and GraphAgent-Reasoner struggles to solve them. The following shows its attempt on the Hamilton Path problem.
@@ 8.1 The GraphInstruct Dataset
8.2 Other Execution Examples
@@
8.3 Distributed Algorithms
Example of distributed algorithms in distributed algorithm library.
Shortest Path. This distributed algorithm computes the shortest distance from a source node to all other nodes in a weighted graph. Each node maintains a state variable distance representing the known shortest distance from the source node. During initialization, the source node sets its distance to 0, while all other nodes set their distance to infinity. In each round, nodes send messages containing new_distance to their neighbors, where new_distance is calculated as the sum of the current node’s distance and the edge weight to each neighbor. Upon receiving messages, each node updates its distance if the received new_distance is smaller than its current distance. The algorithm terminates when no node updates its state, ensuring that all nodes have computed their shortest distances from the source. See Figure 6.
Connectivity. This distributed algorithm determines connected components in a graph by propagating component identifiers. Each node maintains a state variable Component_Id representing the identifier of the connected component it belongs to. During initialization, each node sets its Component_Id to its own node ID. In each round, nodes send messages containing their current Sender_Component_Id to all neighbors. Upon receiving messages, each node compares the received Sender_Component_Id with its current Component_Id. If the received ID is smaller, the node updates its Component_Id to the smaller value; otherwise, it keeps its current Component_Id unchanged. The algorithm terminates when no node updates its state, at which point nodes with the same Component_Id belong to the same connected component. See Figure 7.
Example of distributed algorithms designed by the Master LLM.
PageRank. This distributed algorithm computes the importance score (PageRank) of webpages in a directed graph. Each node maintains a state variable rank representing the importance of the webpage, initialized to 1.0. In each round, nodes calculate their contribution as rank / out_degree, where out_degree is the total number of outgoing links, and send this contribution to all connected neighbors. Upon receiving messages, each node sums all received contribution values and updates its rank using the PageRank formula: new_rank = damping_factor * (sum_of_received_contributions) + (1 - damping_factor) * initial_rank, where damping_factor is typically set to 0.85. The algorithm iterates until the rank values converge, ensuring that each node has computed its final PageRank score. See Figure 8.
Hamilton Path. This distributed algorithm determines whether a Hamiltonian path exists in a graph by propagating path information. Each node maintains two state variables: isInPath (a boolean indicating if the node is part of the Hamiltonian path) and pathLength (an integer representing the length of the path ending at this node). During initialization, a starting node (e.g., Node 0) sets isInPath to True and pathLength to 1, while all other nodes set isInPath to False and pathLength to 1. In each round, nodes with isInPath = True send messages containing their pathLength + 1 and isInPath status to neighbors. Upon receiving messages with isInPath = True, each node updates its pathLength to the maximum of its current value and the received pathLength, and sets isInPath to True if the received path is longer. The algorithm terminates when no updates occur, and a Hamiltonian path exists if any node reaches a pathLength equal to the total number of nodes. See Figure 9.
Subgraph Matching. This distributed algorithm verifies whether there exists a subgraph in graph that is isomorphic to a given subgraph . Each node in graph maintains two state variables: NodeMatch (a list of potential matches for nodes in the subgraph ) and MatchedFlag (a boolean indicating if the node is currently matched to a node in the subgraph). During initialization, each node in sets NodeMatch as an empty list and MatchedFlag as false, while each node in subgraph initializes its unique identifier and sets MatchedFlag as false. In each round, nodes send two types of messages: MatchInfo (containing the node identifier and its list of current potential matches) and EdgeConnection (indicating the directed neighbors that the sending node is connected to). Upon receiving MatchInfo messages, each node checks if the identifier matches any node in subgraph , and if a match is found, updates NodeMatch to include the new match and sets MatchedFlag to true. For EdgeConnection messages, nodes update their list of potential neighbors or confirm connectivity with received nodes, ensuring to track relationships needed for further matching. The algorithm terminates when no new matches are found and the MatchedFlag remains false, retaining the current state while preparing for potential future updates. See Figure 10.